# Deep Learning Reduced Order Modelling on Parametric and Data driven domains

### v1

@@ 3.2 Domains with varying number of parameters
Figure 8: Exact solutions (left) not multiplied by characteristic function, model prediction (middle) not multiplied by characteristic function, relative error (right) multiplied by characteristic function for slightly deformed circles with parameters $\beta=1.19617$, $\theta=1.71657$. Ellipses have centers, $(0.66555,0.35233),(0.56892,0.71108),(0.25424,0.21219)$ and radii of non-deformed circles $0.11191,0.12318,0.10707,0.10653$ respectively. Small half-axes are $20\%$ smaller than large half-axes for each ellipse. Even though the model was not trained on elliptic holes, we can see a small relative error further from the boundary. The error is the largest at the boundary since the solution autoencoder did not capture ellipses in the original training dataset. This demonstrates that the model can work with small perturbations of the boundary.
Conclusion
In this work, we present the extension of DL-ROM to various 2D domains. This approach can be extended to 3D problems. Our algorithm can be used in problems where domains come from measurements. However, this comes at the expense of the main disadvantage of this approach, which is the first step of interpolating into a rectangular domain. In some problems, such as 3D thin pipes with torsion, rectangular data can be much larger than the actual number of nodes required to describe the pipe, which can be a computational bottleneck. In future work, we plan to mitigate this issue as well. We demonstrated that meaningful geometrical parametrization can be obtained using neural networks, which may be useful in other reduced-order model techniques as well.
Acknowledgments
@@
This research was performed using the advanced computing service provided by the University of Zagreb University Computing Centre - SRCE. This research was supported by the Croatian Science Foundation under the project number IP-2022-10-2962 (I.M., B.M. and D.V.).

### v2

@@ 4.2 Domains with varying number of parameters
Figure 9: The exact solution (left) not multiplied by characteristic function, the model prediction (middle) not multiplied by a characteristic function, relative error (right) multiplied by characteristic function for slightly deformed circles with parameters $\beta=9.643$, $\phi=4.129$. Ellipses have centers, $(0.601,0.0481)$, $(0.441,0.0.206)$, $(0.0.789,0.0.206)$, and $(0.334,0.770)$, and radii of non-deformed circles $0.113$, $0.105$, $0.101$, and $0.118$, respectively. Small half-axes are $15\%$ smaller than large half-axes for each ellipse. While relative error is highest near domain boundaries due to unseen shapes during training, the model maintains low error elsewhere, showing robustness to domain perturbation.
Figure 10: Histogram of relative errors on test set for model with ROM encoding dimension 30 and domain encoding dimension 30. Only $38$ out of $1000$ samples have relative error larger than $0.05$.
4.3 Navier-Stokes equations
In this section, we evaluate the DL-DA-ROM framework on the Navier–Stokes equations, a significantly more complex and challenging problem. The Navier–Stokes equations govern fluid flow and include significant nonlinearities, providing a challenging and realistic scenario to test the effectiveness and robustness of our methodology in practical applications.
The considered scenario involves fluid flow through a cavity with a geometrically complex boundary, motivated by the simulations of blood flow through vessels with aneurysms. Specifically, the domain is characterized as a channel with a variable width, representing an aneurysm-like enlargement, and further complicated by the presence of a circular obstacle to introduce additional geometric complexity. We vary the Reynolds number between 1 and 100.
Furthermore, to assess the robustness of our model, we deliberately test parameter configurations that lie outside the original training distribution. This allows us to verify that our model can handle moderate extrapolation scenario, ensuring stability and reasonable accuracy for previously unseen, but not drastically different, geometries and flow conditions. Precisely, we consider the following boundary value problem:
[…]
where the domain $\Omega(\lambda))$ is a cavity without a circle. The length of the domain is $L=4.5$ and $\sigma$ is the fluid stress tensor. Top side is deformed as $R+he^{-w(x-1.25)^{2}}+s(x)$, where the spline $s$ is given by $9$ points $(0,0),(0.25,0),(x_{s1},y_{s1}),(x_{s2},0),(2.2,0),(x,he^{(-w(-L/2)^{2})}-he^{-w(x-L/2)^{2}})$ for $x=3.0,3.5,4.0,4.5$ and the bottom is deformed as $he^{w(x-1.25)^{2}}-s(x)$. The parameters are:
  • undeformed pipe width $R\in[0.1,0.2]$
  • obstacle radius $rCirc\in[0.1,0.15]$
  • $x_{s1}\in[0.5,1]$
  • $y_{s1}\in[0.15,0.25]$
  • $x_{s2}\in[1.5,2]$
  • obstacle center $(x_{0},y_{0})\in[0.5,1.5]\times[0.05,0.1]$
  • inflow amplitude $a\in[5,15]$
  • $w\in[0.05,0.2]$
  • $h\in[0.15,0.2]$
  • $Re\in[1,100]$
In total, we have two PDE parameters: inflow amplitude $a$ and Reynolds number $Re$ and $9$ geometric parameters: pipe width $R$, obstacle radius $rCirc$, obstacle center $(x_{0},y_{0})$, spline points $x_{s1},y_{s1},x_{s2}$ and $w$ and $h$ determining the shape of the cavity. The solutions are interpolated on the $256\times 512$ grid on $0.5,0.725]\times[0,4.5]$. We train the model to predict the velocity in $x$ and $y$-direction. The autoencoder is trained on the velocity components in both the $x$ and $y$-directions, while all reported relative errors correspond to the velocity magnitude. A comparison of model performance using calculated versus exact domain parameters is presented in Table 7, with the representative results shown in Figure 11. We emphasize that despite the increased complexity of both the governing equations and the domain geometry compared to the advection-diffusion case, the average error only slightly increases and remains well within acceptable limits for reduced-order modeling.
test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 exact domain parameters only 0.0412 0.0412 0.0335 0.0335 calculated domain parameters 0.0361 0.0348 0.0379 0.0333
Table 7: Test relative errors obtained using different ROM and the domain encoding dimensions. We differentiate between cases where geometric parameters are exactly known (1st row) and when we use only geometric parameters obtained by the autoencoder (2nd row). We can see that the use of the autoencoder domain parameterization does not negatively affect DL-ROM performance. We report on the best model for each ROM encoding dimension and domain encoding dimension. Most Models trained on calculated domain parameters perform better than the models trained using only exact domain parameters. Even though results are similar, larger fully-connected neural network is needed in case of calculated domain parameters (4 layers with 2048 neurons vs 4 layers with 512 neurons).
Figure 11: The exact solutions (top left), model prediction (top right) and relative error (bottom) for model trained on $1000$ random parameter samples, with 2020-dimensional domain parametrization obtain using $a=10.34$ and $Re=30.1$.
We further evaluate the sensitivity of the model to domain perturbations by testing parameter configurations that fall outside the range of the original training set. Specifically, we consider domains with wider cavities ($w\in[0.2,0.25]$), using the model with ROM and domain encoding dimensions both set to $30$. Additionally, we increase the Reynolds number slightly beyond the training range, with $Re\in[100,110]$, while other parameters remain within the original bounds. A total of $100$ such samples were generated, resulting in an average relative error of approximately $0.065$.
We closely examine outlier cases and find that one source of increased error occurs when the domain autoencoder fails to accurately reconstruct the domain geometry (see Figure 12). Nevertheless, even in such cases, the method provides a reasonable approximation with a relative error of $0.138$. Conversely, Figure 13 illustrates a case where the domain encoding performed well, resulting in a relative error of $0.087$. These examples indicate that, despite some reconstruction issues, the model remains robust and produces acceptable solutions even under moderate parameter and geometry shifts.
Figure 12: Out of distribution sample with the highest relative error of $0.138$. Domain reconstruction (top left), correct solution (top right), prediction (bottom left), relative error (bottom right). Domain autoencoder is not able to properly reconstruct the domain, producing artificial holes in the domain. This implies that calculated domain parametrization is not appropriate for this domain. However, the model still finds a reasonable solution.
Figure 13: Typical out of distribution sample. Domain reconstruction (top left), correct solution (top right), prediction (bottom left), relative error (bottom right). In this case, the domain autoencoder is able to reconstruct the domain, thus able to extract more meaningful representation.
The distribution of the relative error on the test set for the best model reported in Table 7 is shown in Figure 14. We can see that the model predictions are stable across the test dataset.
Figure 14: Histogram of relative errors on test set for the model with ROM encoding dimension 30 trained with exact domain parameters (blue), and trained with domain encoding dimension 30 (orange).
5 Conclusion
We presented a method for building reduced-order models for PDEs on a parameter-dependent domains, including scenarios where the domain geometry is not explicitly given but instead defined through images or measurements. Unlike many existing ROM techniques that require consistent meshing or graph-based representations, our approach directly utilizes bitmap images of domain characteristic functions. This makes it particularly suitable for practical applications where mesh generation is costly, unreliable, or unavailable—such as simulations derived from imaging data or experimental measurements.
Our results demonstrate that the learned domain encoding is not only compact but also meaningful; in cases with known geometric parameters, our encoding achieved similar accuracy to models using exact parameterizations. Furthermore, our method effectively handled significant variations in domain shape, changes in the number of features (such as holes), and moderate boundary perturbations without losing stability. Although the use of domain parametrization obtained via the CNN autoencoder successfully enables mapping from physical parameters to the latent solution space, we observed that this configuration generally requires a larger network capacity to achieve comparable accuracy. This behavior likely stems from a partial misalignment between the domain parametrization and the reduced-order latent space, which increases the effective nonlinearity of the mapping. Future work will focus on improving the consistency between the autoencoder latent representation and the parameter space, for example through joint or coupled training strategies or latent regularization terms. Such approaches could reduce the required network size and improve generalization while preserving the interpretability and robustness of the proposed DA-DL-ROM framework. Conceptually, our approach extends naturally to three-dimensional problems. However, from a practical perspective, computational efficiency remains an important challenge due to the increased size of the input grids. Improving computational efficiency for three-dimensional implementations is currently a focus of our ongoing work.
Additionally, the learned domain representations may offer utility beyond DL-ROM frameworks, potentially benefiting other situations requiring domain parameterization where explicit parameters are unavailable or in different machine learning-based model reduction contexts.
6 Acknowledgments
@@
This research was carried out using the advanced computing service provided by the University of Zagreb University Computing Centre - SRCE. This research was supported by the Croatian Science Foundation under the project number IP-2022-10-2962 (I.M., B.M. and D.V.). M.B. has been partly supported by the National Science Foundation via grants NSF DMS-2208219 and NSF DMS-2205695.